OpenStudy (anonymous):

HOW DOES THETA VARIES IN PARAMETRIC EQUATION OF SPHERE....LIKE IN UPPER HEMISPHERE,RIGHT HEMISPHERE

OpenStudy (anonymous):

\[the equation of sphere is x^2+y^2+z^2=a^2 its parametric eqn is x=\rho \sin \phi \cos \theta,y=\rho \sin \phi \sin \theta,z=\rho \cos \phi\] \[how does \theta varies\]